OpenStudy (kiiraa_x3):

f(x)=1(x-12)^2+9 ... what are the steps to putting this in standard form? HELP!

OpenStudy (kiiraa_x3):

@dan815

OpenStudy (kiiraa_x3):

@nincompoop @nuccioreggie @jagr2713 @King.Void.

OpenStudy (kiiraa_x3):

@fierstar123 @AriPotta @@SolomonZelman @iGreen @mathstudent55 @mathmath333 @matt101

OpenStudy (matt101):

You just need to expand and simplify. First square the stuff in the brackets. Then collect all like terms. Your equation will then be in standard form!

OpenStudy (kiiraa_x3):

Can you like solve it and graph it for me please so I can understand it better? @matt101

OpenStudy (matt101):

First square the brackets. If you have (a+b)^2 and you expand it, it works out to a^2 + 2ab + b^2. In this case, a=x and b=-12. So, if we have (x-12)^2, that equals x^2+2x(-12)+(-12)^2, which is x^2-24x+144. Now let's look at our equation: \[f(x)=(x-12)^2+9\]\[f(x)=(x^2-24x+144)+9\]\[f(x)=x^2-24x+153\]

OpenStudy (matt101):

That's your equation in standard form. Standard form means your equation is in the form of f(x)=Ax^2+Bx+C, where A, B, and C are constants. The form your equation started in is called vertex form, because you can easily see the coordinates of the vertex of the parabola in the equation.

OpenStudy (kiiraa_x3):

thanks so much :) @matt101

OpenStudy (kiiraa_x3):

@Mashy How do you graph this?

OpenStudy (kiiraa_x3):

@mathhead247

OpenStudy (kiiraa_x3):

@Abhisar @CGGURUMANJUNATH @Michele_Laino how to find the average rate of change for f(x) from x = 4 to x = 7.

OpenStudy (kiiraa_x3):

@oblivion20

OpenStudy (kiiraa_x3):

@SneliS @SolomonZelman